A proton moves through a uniform magnetic field of magnitude 2.0 μT = 2.0


x 10^-6 T at a speed of 5.00 x 10^6 m/s.



(a) What is the magnitude of maximum


magnetic force it can experience?



(b) If the proton is replaced by an electron,


will it experience the same maximum force?

1
Expert's answer
2022-04-06T10:13:24-0400

(a) What is the magnitude of maximum magnetic force it can experience?

"F=qvB\\
=1.6\times 10^{-19}\times 5.00 \times10^6\times 2.0\times10^{-6}\\
=1.6\times 10^{-18}\rm \: N"

(b) If the proton is replaced by an electron, will it experience the same maximum force?

The charge of an electron is the same as a proton, so the force will be the same.